OpenStudy (nastech):

please how do I find the sum from k=1 to infinity of 3k/(k+4)

OpenStudy (anonymous):

Divide everything by k

OpenStudy (klimenkov):

\[\sum_{k=1}^{\infty}\frac{3k}{k+4}\]Try to find a limit\[\lim_{k\rightarrow\infty}\frac{3k}{k+4}\]

OpenStudy (anonymous):

are you looking for the sum??? the sum is unbounded as k goes to infinity...

OpenStudy (anonymous):

3k/(k+4) =3(k+4-4)/(k+4)= 3(1-4/(k+4)) if k tends to infinity 4/k+4 tends to 0 hence the required sum is 3

OpenStudy (anonymous):

the limit approaches to 3

OpenStudy (anonymous):

so is this what you want? \(\large \sum_{k=1}^{\infty}\frac{3k}{k+4} \)

OpenStudy (klimenkov):

\[\lim_{k\rightarrow\infty}a_k\ne0\Rightarrow\sum_{k=1}^\infty a_k=\infty\]

OpenStudy (anonymous):

Why are you guys changing the question? The asker said "sum from k=1 to infinity of 3k/(k+4)" Isn't this what he want: \(\Large \sum_{k=1}^{\infty}\frac{3k}{k+4} \) ??? and not the limit?

OpenStudy (nastech):

finding the limit will enable us to determine whether it converges or diverges

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty}(\frac{ 3k }{ k+4 })=\infty\]

OpenStudy (anonymous):

go for it then....

OpenStudy (anonymous):

Like I said earlier, the sum is unbounded..... http://www.wolframalpha.com/input/?i=sum+%283*k%29%2F%28k%2B4%29+from+k%3D1+to+infinity

OpenStudy (anonymous):

Ooh, I screwed that up, sorry.....

OpenStudy (turingtest):

as is stated above, the limit of the sequence (the summand) must converge to 0 or else the sum is undefined. note: just because the limit of the summand goes to zero does *not* mean the series converges, just that it might.

OpenStudy (turingtest):

limit as n->infty that is...

OpenStudy (turingtest):

since the limit of the summand is not zero, we know the series does not converge, as was stated earlier

OpenStudy (turingtest):

there are many other ways to see that this series does not converge btw

OpenStudy (nastech):

I'm getting confused

OpenStudy (anonymous):

@Nastech just check the initial terms for k=1 , a1 = 3/4 , k=2, a2 =6/5 and for k=3 a3=9/7....if u see them clearly the series is always increasing and will reach to infinity for the infiniteth term...i.e summation of this series is unbounded or infinity

OpenStudy (turingtest):

look, there are a number of tests for convergence I think you should acquaint yourself with some the one I think best for this series is the limit test

OpenStudy (anonymous):

@Nastech do u get it now

OpenStudy (turingtest):

the limit test is that for a sum\[\sum_{n=k}^\infty a_n\]if\lim_{n\to\infty}a_n\neq0\]then the series diverges, if \[\lim_{n\to\infty}a_n=0\]then it \(might\) converge, and we need another test to determine (ratio test, integral test, etc.)

OpenStudy (turingtest):

edit* if\[\lim_{n\to\infty}a_n\neq0\]then the series diverges, if

OpenStudy (nastech):

thanks a lot I get it now

OpenStudy (nastech):



OpenStudy (turingtest):

are you posting a totally separate DE ? please post each question separately, thanks.